# A systematically-selected sample of luminous, long-duration, ambiguous nuclear transients

### v1

@@ 4 Comparisons between the variability of ANTs and AGN
We assess the variability of the 10 ANTs in the main sample, the 59 ‘gold’ candidates returned by our linear-decline, high-amplitude selection as well as the ‘faster’ and ‘bumpier’ samples. We also compute the SF for all ZTF nuclear transient candidates detected in the first year of ZTF, most of which are likely AGN. We restrict to these candidates so that we have at least a 5 yr baseline. We fit the difference-imaging light curves with a DRW and determine their and . To do so we fit the observer-frame ZTF -band light curves using the taufit202020https://github.com/burke86/taufit package. The results are shown in Fig. 8: SDSS AGN (MacLeod et al., 2010) and ZTF nuclear candidates are shown in blue and dark red contours, respectively while the SFs for ANTs in our gold sample are shown in open gold diamonds, and the SFs for ANTs are shown in open green hexagons. In general, DRW fits are unreliable if the observed baseline is not significantly longer than . For the 7229 ZTF nuclear light curves, the maximum possible baseline is d, indicated by a horizontal line in the figure. Results above this line are unreliable while results below the line could be lower limits since longer baselines were not measureable. The ZTF candidates are all shifted to higher amplitude compared to the SDSS AGN, likely because all of these events are transient candidates detected in difference imaging and thus must be variable above the ZTF detection limit in the difference image (the threshold at is about 0.6 mag.
@@ 4.1 Rise time versus amplitude
The 59 ‘gold’ ANT candidates are selected to be smoothly evolving. Although the DRW is still a valid description of a low-amplitude, long timescale variability it is far from an ideal model for such transient variability. Nevertheless, the resulting all lie to the upper edge of the contour from ZTF, and are (by selection) larger than almost all MacLeod et al. (2010) AGN. The main ANT sample is clearly shifted compared to the full ‘gold’ sample which includes objects with shorter and smaller . To assess the significance of the largest flare, we measure the observer-frame rise time from first detection to peak brightness of the -band light curve (not -corrected) and plot these against the change in difference magnitude over the same timescale. These are shown as filled points in Fig. 8, for the main sample of 10 plus the ‘gold’, ‘bumpier’ and ‘faster’ selections. These measurements clearly indicate higher amplitude variability than measured by the DRW, and indicate a shorter timescale for that variability. There is a clear preference for the ANTs to have shorter rises from first detection to peak brightness compared to the ‘gold’, ‘bumpier’ and ‘faster’ samples. This is despite the ‘faster’ sample being selected for having a shorter decline, and indicates that most events picked up by that selection have slow rises and fast declines, indicative of stochastic variability rather than transient behaviour.
We construct a region in rise-time vs rise amplitude space that includes all 10 ANTs from the main sample. This region is bounded by (equivalent to 1.6 mag) and (equivalent to d). We suggest that any light curves falling outside this range are likely due to stochastic AGN variability. Within this region are 9 ‘platinum’ events that are not in our main sample. Of these, one is a peculiar TDE, two are SNe in a star-forming knots, and three are ANTs that pass our visual inspection but do not have spectra. Thus there are only 4 true AGN ‘contaminants’ when applying these selection criteria and we advocate their use in future surveys.
@@ 5 Discussion
We have searched for analogues of the extremely luminous transient AT2021lwx, and have systematically identified a futher sample of slow and smoothly evolving nuclear transients in the ZTF data stream. Of these transients, only one (ZTF19aailpwl/AT2019brs) has been published as a transient of interest. Another, ZTF21abxowzx/AT2021yzu, was publicly classified as an AGN. With this pipeline we identify four further events with photometric and spectroscopic properties consistent with ANTs. ZTF18aczpgwm/AT2019kn has broad, high equivalent-width emission lines as well as very strong iron complex. ZTF19aamrjar is the longest-lived of the sample and has the highest integrated radiated energy. Its spectrum resembles a quasar, and it has red-winged asymmetric Balmer lines, yet its photometric evolution is entirely smooth for over 2 years in the rest frame. ZTF20aaqtncr/AT2021fez is a lower amplitude flare whose emission lines resemble a type II AGN. ZTF20abodaps/AT2020afep rises on the timescale of a regular SN or TDE, but declines far, far slower. All of these transients display a MIR flare consistent with circumnuclear dust echoes.
@@ 5.1 TDEs, CLAGNs, or something in between?
The samples presented in this work add to a growing list of ANTs discovered in untargetted searches. Some previously published events (e.g. PS1-10adi, PS16dtm, Gaia16aaw, ASASSN-17jz, Gaia18cdj) have similar smooth rises and declines to the ZTF sample. On the other hand, another set of events show differing light curves with either fast rises, bumpy declines, or plateaus (e.g. ASASSN-15lh, OGLE17aaj, AT2017bgt, ASASSN-18jd, Swift J221951-484240). Nevertheless, all but ASASSN-15lh are spectroscopically similar, dominated by strong narrow Balmer emission lines (and ASASSN-15lh has been claimed to be a TDE by Leloudas et al. 2016). These observations indicate that slow-moving gas is nearly ubiquitous in ANTs but that the geometry and/or central energy source may not be, leading to the different light curve shapes.
@@ 5.1.1 Extreme instability in an existing accretion disk
Arguably the most simple explanation of ANTs is an intrinsic change in the structure of an existing accretion disk (e.g. Sniegowska et al., 2020). The existence of the disk before the flare is required in many ANTs which are found in NLSy1 galaxies or other AGN. This scenario is comprehensively addressed in Cannizzaro et al. (2020) where it is shown that the rise timescales of ANTs (months) are far too short compared to the dynamical timescale of a classical accretion disk (>70 years).
@@ 5.1.3 Tidal disruption event of a massive star
Here we fit the light curves of the main sample with the Modular Open Source Fitter for Transients (MOSFiT Guillochon et al., 2018) using the TDE model of Mockler et al. (2019). We achieve reasonable fits to all the light curves except ZTF20abodaps/AT2020afep, for which the rise time is too fast compared to the decline time. The best fit stellar and black hole masses are presented in Table 7. The most eyecatching is ZTF19aamrjar, which requires and . Such a scenario is far beyond the scope of the MOSFiT models which derive from simulations of solar mass stars. All of the others have parameters more in line with what might be expected for extreme but plausible TDEs: (with the exception of ZTF18aczpgwm/AT2019kn for which ) and .
@@ 5.2 Volumetric rate of ANTs
The intrinsic rate of transient events provides constraints on the physical mechanisms and plausible scenarios causing them. For example, for ANTs to be caused by high stellar mass TDEs, the rates should agree with the expected population of high-mass stars in SMBH loss cones. Such rate measurements are important for the inclusion of ANTs in simulations of galaxies and their nuclei, the understanding of stellar dynamics and black hole mass distributions (e.g. Stone & Metzger, 2016) as well as for explaining the chemical abundances of AGN (e.g. Kochanek, 2016). We estimate the volumetric rate of ANTs by using the seven events in the main sample that were selected by our pipeline. We limit the calculation to , the highest redshift event (ZTF20abrbeie/AT2021lwx). For each event we calculate the maximum redshift that it would have passed our cuts, using the black body properties from Table 4. Using that maximum redshift, we then weight each transient by the maximum observable volume, (Schmidt, 1968). We assume the sky area of the public survey to be deg2 (Ho et al., 2023), and the survey time to be 4.77 yr (we cut the selection at 31/12/2023/MJD 60309 so for the transient to last 1 yr, it must be first detected by MJD 59943). Counting the seven events in this manner, we estimate a lower limit on the rate Mpc-1 yr-1. Removing ZTF20aaqtncr/AT2021fez, which we classified as Unclear, gives a rate of Mpc-1 yr-1. We note that these measurements are times greater than those for ‘extreme’ ANTs (ENTs) measured by Hinkle et al. (2024), a sample of three with which we share one common event (ZTF20abrbeie/AT2021lwx). These rates are consistent with a direct extrapolation of the measured TDE luminosity function (i.e. rate per unit peak luminosity) in ZTF measured by Yao et al. (2023a). However, above the Hills mass of for a solar mass star the TDE rate should be heavily suppressed (e.g. Stone et al., 2020), and indeed such a suppression is observed (van Velzen et al., 2019; Mockler et al., 2019; Nicholl et al., 2022). Wiseman et al. (2023) estimated that the rate of TDEs from stars should be that of solar mass stars based on the slope of the initial mass function (IMF) and the lifetime of massive stars. Applying this fraction to the observed volumetric TDE rate of Mpc-1 yr-1 (Yao et al., 2023a) we might expect Mpc-1 yr-1 from stars. The TDE model fits to our ANT light curves returned stellar masses between , with seven in the range . Based upon the arguments of lifetimes and the IMF would expect significantly more TDEs from stars in this mass range than . The observed rate Mpc-1 yr-1 is thus consistent with the expectation for TDEs of stars in the mass range. The increased depth of the Vera C. Rubin Observatory’s Legacy Survey of Space and Time (LSST Ivezic et al., 2019) will enhance the rate of observed TDEs Bučar Bricman et al. (2023) while hundreds will be spectroscopically classified by the Time Domain Extragalactic Survey (TiDES; Swann 2020, Frohmaier et al., in prep) on the 4-metre Multi-Object Spectrograph Telescope. These observations will allow a detailed measurement of the rate of massive-star TDEs and ANTs.
@@ 5.3 Uniqueness of AT2021lwx
Despite a thorough search, no events were found with observed properties that are comparable properties to AT2021lwx. Firstly, every event discovered using the linear fit was found with a clear host galaxy, with AT2021lwx itself the only hostless event passing the cuts. No other event in our sample has a luminosity as low as the limit of mag for the host of AT2021lwx. Assuming a black body SED for the transient, two events (ZTF19aamrjar and ZTF20abodaps/AT2020afep) from the main sample come within dex of the peak UV-optical pesudo-bologmetric luminosity of AT2021lwx (=45.85). The light curve of ZTF19aamrjar is the most similar to AT2021lwx, and since it evolves much more slowly, the integrated energy release is larger than AT2021lwx. The spectrum is however vastly different: the Balmer line profiles are broad and asymmetric, and there is strong narrow oxygen. ZTF20abodaps/AT2020afep has a similar spectrum, with no oxygen present. However, the sharp rise time is inconsistent with AT2021lwx.
@@ 6 Conclusions
We find six transients with long-duration ( yr), smooth, luminous ( erg s-1) light curves similar to ZTF20abrbeie/AT2021lwx, five of which we classify as ANTs and one as uncertain. Our method also recovers ZTF20abrbeie/AT2021lwx itself.
@@
If these ANTs, along with those compiled from other surveys, are indeed TDEs then they emission mechanism for the bulk of the radiation must be different to the canonical TDE of a star around a black hole in order to produce the markedly different optical spectra. The high dust covering fractions imply that ANTs may be heavily-obscured TDEs for which we observe reprocessed, slow-moving material.
@@ Acknowledgements
We thank S. Smartt and the Pan-STARRS team for providing data points of ZTF19aamrjar, ZTF20abodaps, ZTF18aczpgwm, ZTF19aatubsj, ZTF20aanxcpf to support the figures in the appendix. We are grateful to Matt Stepney and Daniel Kynoch for their advice and insights on AGN spectra and Sandra Raimundo for discussions about CLAGNs.
@@ A.2 Faster light curves
The second variation is to relax the light curve duration to yr. We also relax the number of detections to 25. This allows for faster-evolving events, and events with limited visibility (e.g. a 150 d rise was unobservable from Palomar but a subsequent 200 d decline was observed). It also means the contamination from regular transients is increased. 95 transients pass this ‘faster’ selection, of which 11 pass visual inspection having not already been included in the gold sample. Of the 11, there are three spectroscopically classified SNe IIn: ZTF19abpvbzf appears in a knot of star formation in an irregular galaxy, ZTF22aadesjc appears visually off-centre in a dwarf galaxy, but ZTF22aaurfwa=SN2022hem is only classified by a low-resolution spectrum and may be an ANT. Four objects are spectroscopically classified TDEs: ZTF18abxftqm=AT2018hco (van Velzen et al., 2018), ZTF20abnorit=AT2020ysg (Yao et al., 2023a), ZTF20achpcvt =AT2020vwl (Hammerstein et al., 2021). ZTF23aadcbay=AT2023cvb (Yao et al., 2023b). ZTF18aaqkcso is a type Ia SN. In Fig 9 we show the light curve of ZTF20aaurfwa/SN2022hem, along with one ZTF20abnorit/AT2020ysg which is classified as TDE-featureless and thus is ambiguous. ZTF20abgxlut has an ANT-like spectrum212121https://www.wiserep.org/object/15336 with narrow Balmer emission along with He ii, and the lightcurve is only marginally shorter than our 1 yr threshold.

### v2

@@ 4 Comparisons between the variability of ANTs and AGN
We assess the variability of the 11 ANTs in the analysis sample, the 59 ‘gold’ candidates returned by our linear-decline, high-amplitude selection as well as the ‘faster’ and ‘bumpier’ samples. We also compute the SF for all ZTF nuclear transient candidates detected in the first year of ZTF, most of which are likely AGN. We restrict to these candidates so that we have at least a 5 yr baseline. Instead of the difference-imaging magnitudes, which amplify differences with respect to some baseline, we obtain photometry from the ZTF Public Data Release 22222222https://www.ztf.caltech.edu/ztf-public-releases.html (Masci et al., 2018) that includes any baseline AGN luminosity as well as the host galaxy contribution and can be compared directly to the well-studied SDSS sample. We fit the observer-frame -band light curves with a DRW using the taufit232323https://github.com/burke86/taufit package, fitting for the and parameters. In an ideal case we would use rest frame light curves, but the majority of ZTF candidates do not have measured redshifts to allow for a -correction. The results are shown in Fig. 9: SDSS AGN (MacLeod et al., 2010) and ZTF nuclear candidates are shown in blue and dark red contours, respectively while the SFs for ANTs in our gold sample are shown in open gold diamonds, and the SFs for ANTs are shown in open green hexagons. In general, DRW fits are unreliable if the observed baseline is not significantly longer than . For the 7229 ZTF nuclear light curves, the maximum possible baseline is d, indicated by a horizontal line in the figure. Results above this line are unreliable while results below the line could be lower limits since longer baselines were not measureable. The ZTF candidates are all shifted to higher amplitude compared to the SDSS AGN, likely because all of these events are transient candidates detected in difference imaging and thus must be variable above the ZTF detection limit in the difference image (the threshold at is about 0.6 mag.
@@ 4.1 Rise time versus amplitude
The 59 ‘gold’ ANT candidates are selected to be smoothly evolving. Although the DRW is still a valid description of a low-amplitude, long timescale variability it is far from an ideal model for such transient variability. Nevertheless, the resulting all lie to the upper edge of the contour from ZTF, and are (by selection) larger than almost all MacLeod et al. (2010) AGN. The analysis ANT sample is clearly shifted compared to the full ‘gold’ sample which includes objects with shorter and smaller . To assess the significance of the largest flare, we measure the observer-frame rise time from first detection to peak brightness of the -band light curve (not -corrected) and plot these against the change in difference magnitude over the same timescale. These are shown as filled points in Fig. 9, for the analysis sample of 11 plus the ‘gold’, ‘bumpier’ and ‘faster’ selections. These measurements clearly indicate higher amplitude variability than measured by the DRW, and indicate a shorter timescale for that variability. There is a clear preference for the ANTs to have shorter rises from first detection to peak brightness compared to the ‘gold’, ‘bumpier’ and ‘faster’ samples. This is despite the ‘faster’ sample being selected for having a shorter decline, and indicates that most events picked up by that selection have slow rises and fast declines, indicative of stochastic variability rather than transient behaviour.
We construct a region in rise-time vs rise amplitude space that includes all 11 ANTs from the analysis sample. This region is bounded by (equivalent to 1.6 mag) and (equivalent to d). We suggest that any light curves falling outside this range are likely due to stochastic AGN variability. Within this region are 9 ‘platinum’ events that are not in our analysis sample. Of these, one is a peculiar TDE, two are SNe in a star-forming knots, and three are ANTs that pass our visual inspection but do not have spectra. Thus there are only 4 true AGN ‘contaminants’ when applying these selection criteria and we advocate their use in future surveys.
@@ 5 Discussion
We have searched for analogues of the extremely luminous transient AT2021lwx, and have systematically identified a further sample of slow and smoothly evolving nuclear transients in the ZTF data stream. Of these transients, only one (ZTF19aailpwl/AT2019brs) has been published as a transient of interest. Another, ZTF21abxowzx/AT2021yzu, was publicly classified as an AGN. With this pipeline we identify four further events with photometric and spectroscopic properties consistent with ANTs. ZTF18aczpgwm/AT2019kn has broad, high equivalent-width emission lines as well as very strong iron complex. ZTF19aamrjar is the longest-lived of the sample and has the highest integrated radiated energy. Its spectrum resembles a quasar, and it has red-winged asymmetric Balmer lines, yet its photometric evolution is entirely smooth for over 2 years in the rest frame. ZTF20aaqtncr/AT2021fez is a lower amplitude flare whose emission lines resemble a type II AGN. ZTF20abodaps/AT2020afep rises on the timescale of a regular SN or TDE, but declines far, far slower. All of these transients display a MIR flare consistent with circumnuclear dust echoes.
@@ 5.1 TDEs, CLAGNs, or something in between?
The samples presented in this work add to a growing list of ANTs discovered in untargeted searches. Some previously published events (e.g. PS1-10adi, PS16dtm, Gaia16aaw, ASASSN-17jz, Gaia18cdj) have similar smooth rises and declines to the ZTF sample. On the other hand, another set of events show differing light curves with either fast rises, bumpy declines, or plateaus (e.g. ASASSN-15lh, OGLE17aaj, AT2017bgt, ASASSN-18jd, Swift J221951-484240). Nevertheless, all but ASASSN-15lh are spectroscopically similar, dominated by strong narrow Balmer emission lines (and ASASSN-15lh has been claimed to be a TDE by Leloudas et al. 2016). These observations indicate that slow-moving gas is nearly ubiquitous in ANTs but that the geometry and/or central energy source may not be, leading to the different light curve shapes.
@@ 5.1.1 Extreme instability in an existing accretion disk
Arguably the most simple explanation of ANTs is an intrinsic change in the structure of an existing accretion disk (e.g. Sniegowska et al., 2020). The existence of the disk before the flare is required in many ANTs which are found in NLSy1 galaxies or other AGN. This scenario is comprehensively addressed in Cannizzaro et al. (2020) where it is shown that the rise timescales of ANTs (months) are far too short compared to the dynamical timescale of a classical accretion disk (>70 years), although see Lipunova et al. (2024) for possible origins of flares due to thermal fluctuations in disks.
@@ 5.1.3 Tidal disruption event of a massive star
Here we fit the light curves of the analysis sample with the Modular Open Source Fitter for Transients (MOSFiT Guillochon et al., 2018) using the TDE model of Mockler et al. (2019). We achieve reasonable fits to all the light curves except ZTF20abodaps/AT2020afep, for which the rise time is too fast compared to the decline time. The best fit stellar and black hole masses are presented in Table 6. The most eyecatching is ZTF19aamrjar, which requires and . Such a scenario is far beyond the scope of the MOSFiT models which derive from simulations of solar mass stars. All of the others have parameters more in line with what might be expected for extreme but plausible TDEs: (with the exception of ZTF18aczpgwm/AT2019kn for which ) and .
The mean black hole mass of our sample is with a standard deviation of 0.6 dex. We compare these masses to those of 32 TDEs fit using the MOSFiT TDE model in Nicholl et al. (2022), whose highest black hole mass is with a mean . We perform a two-sample Kolmogorov-Smirnov (KS) test and conclude with a -value of 0.0001 (i.e. 99.9%, or ) that the ANT black hole masses are not drawn from the population of TDE black hole masses. We also compare to the 33 black hole masses of TDEs in Yao et al. (2023a) measured using the host galaxy bulge–black hole mass relation. Their mean mass is and although they extend to higher masses than the TDE black hole masses from MOSFiT, the KS test indicates the ANT population is different at significance.
Given that a significant number of the ANTs occurred in galaxies with likely non-negligible pre-existing accretion, we do not attempt to fit the host galaxy photometry to obtain a stellar mass to attempt to infer black hole masses.
@@ 5.2 Volumetric rate of ANTs
The intrinsic rate of transient events provides constraints on the physical mechanisms and plausible scenarios causing them. For example, for ANTs to be caused by high stellar mass TDEs, the rates should agree with the expected population of high-mass stars in SMBH loss cones. Such rate measurements are important for the inclusion of ANTs in simulations of galaxies and their nuclei, the understanding of stellar dynamics and black hole mass distributions (e.g. Stone & Metzger, 2016) as well as for explaining the chemical abundances of AGN (e.g. Kochanek, 2016). We estimate the volumetric rate of ANTs by using the seven events in the analysis sample that were selected by our pipeline. We limit the calculation to , the highest redshift event (ZTF20abrbeie/AT2021lwx). For each event we calculate the maximum redshift that it would have passed our cuts, using the black body properties from Table 4. Using that maximum redshift, we then weight each transient by the maximum observable volume, (Schmidt, 1968). We assume the sky area of the public survey to be deg2 (Ho et al., 2023), and the survey time to be 4.77 yr (we cut the selection at 31/12/2023/MJD 60309 so for the transient to last 1 yr, it must be first detected by MJD 59943). Counting the seven events in this manner, we estimate a lower limit on the rate Mpc-1 yr-1. Removing ZTF20aaqtncr/AT2021fez, which we classified as Unclear, gives a rate of Mpc-1 yr-1. We note that these measurements are times greater than those for ‘extreme’ ANTs (ENTs) measured by Hinkle et al. (2024), a sample of three with which we share one common event (ZTF20abrbeie/AT2021lwx). These rates are consistent with a direct extrapolation of the measured TDE luminosity function (i.e. rate per unit peak luminosity) in ZTF measured by Yao et al. (2023a). However, above the Hills mass of for a solar mass star the TDE rate should be heavily suppressed (e.g. Stone et al., 2020), and indeed such a suppression is observed (van Velzen et al., 2019; Mockler et al., 2019; Nicholl et al., 2022). Wiseman et al. (2023) estimated that the rate of TDEs from stars should be that of solar mass stars based on the slope of the initial mass function (IMF) and the lifetime of massive stars. Applying this fraction to the observed volumetric TDE rate of Mpc-1 yr-1 (Yao et al., 2023a) we might expect Mpc-1 yr-1 from stars. The TDE model fits to our ANT light curves returned stellar masses between , with seven in the range . Based upon the arguments of lifetimes and the IMF would expect significantly more TDEs from stars in this mass range than . The observed rate Mpc-1 yr-1 is thus consistent with the expectation for TDEs of stars in the mass range. The increased depth of the Vera C. Rubin Observatory’s Legacy Survey of Space and Time (LSST Ivezic et al., 2019) will enhance the rate of observed TDEs Bučar Bricman et al. (2023) while hundreds will be spectroscopically classified by the Time Domain Extragalactic Survey (TiDES; Frohmaier et al. 2025) on the 4-metre Multi-Object Spectrograph Telescope. These observations will allow a detailed measurement of the rate of massive-star TDEs and ANTs.
@@ 5.3 Uniqueness of AT2021lwx
Despite a thorough search, no events were found with observed properties that are comparable properties to AT2021lwx. Firstly, every event discovered using the linear fit was found with a clear host galaxy, with AT2021lwx itself the only hostless event passing the cuts. No other event in our sample has a luminosity as low as the limit of mag for the host of AT2021lwx. Assuming a black body SED for the transient, two events (ZTF19aamrjar and ZTF20abodaps/AT2020afep) from the analysis sample come within dex of the peak UV-optical pesudo-bologmetric luminosity of AT2021lwx (=45.85). The light curve of ZTF19aamrjar is the most similar to AT2021lwx, and since it evolves much more slowly, the integrated energy release is larger than AT2021lwx. The spectrum is however vastly different: the Balmer line profiles are broad and asymmetric, and there is strong narrow oxygen. ZTF20abodaps/AT2020afep has a similar spectrum, with no oxygen present. However, the sharp rise time is inconsistent with AT2021lwx.
@@ 6 Conclusions
We find eight nuclear events with long-duration ( d), smooth, luminous ( erg s-1) light curves similar to ZTF20abrbeie/AT2021lwx including ZTF20abrbeie/AT2021lwx itself.
@@
If these ANTs, along with those compiled from other surveys, are indeed TDEs then the emission mechanism for the bulk of the radiation must be different to the canonical TDE of a star around a black hole in order to produce the markedly different optical spectra.
@@ Acknowledgements
We thank the referee for their thoughtful comments which have greatly improved the work. We thank Karri Koljonen for providing the spectra and reduced UVOT data for ZTF22aadesap/AT2022fpx. We thank S. Smartt and the Pan-STARRS team for providing data points of ZTF19aamrjar, ZTF20abodaps, ZTF18aczpgwm, ZTF19aatubsj, ZTF20aanxcpf to support the figures in the appendix. We are grateful to Matt Stepney and Daniel Kynoch for their advice and insights on AGN spectra and Sandra Raimundo for discussions about CLAGNs.
@@ A.2 Faster light curves
The second variation is to relax the light curve duration to yr. We also relax the number of detections to 25. This allows for faster-evolving events, and events with limited visibility (e.g. a 150 d rise was unobservable from Palomar but a subsequent 200 d decline was observed). It also means the contamination from regular transients is increased. 95 transients pass this ‘faster’ selection, of which 11 pass visual inspection having not already been included in the gold sample. Of the 11, there are three spectroscopically classified SNe IIn: ZTF19abpvbzf appears in a knot of star formation in an irregular galaxy, ZTF22aadesjc appears visually off-centre in a dwarf galaxy, but ZTF22aaurfwa=SN2022hem is only classified by a low-resolution spectrum and may be an ANT although the post-peak colour evolution is more consistent with the SN II classification. Four objects are spectroscopically classified TDEs: ZTF18abxftqm=AT2018hco (van Velzen et al., 2018), ZTF20abnorit=AT2020ysg (Hammerstein et al., 2023), ZTF20achpcvt =AT2020vwl (Hammerstein et al., 2021). ZTF23aadcbay=AT2023cvb (Yao et al., 2023b). ZTF18aaqkcso is a type Ia SN. In Fig 10 we show the light curve of ZTF20aaurfwa/SN2022hem, along with ZTF20abnorit/AT2020ysg which is classified as TDE-featureless and thus is ambiguous.